OpenStudy (anonymous):

What is the solution set of the inequality 5 − lx + 4l ≤−3

OpenStudy (anonymous):

5 - |x+4| \(\le\) -3 -|x+4| \(\le\) -3 -5 -|x+4| \(\le\) -8 |x+4| \(\ge\) 8 now solve for x when x+4 \(\ge\) 8 and x + 4 \(\le\) -8